# Practical Use Cases of Neutral Atoms Quantum Computers

### v1

@@ 2 Quantum Computing with neutral atoms
A neutral atom Quantum Processing Unit (QPU) works by trapping and cooling atoms, which typically belong to the alkali group (such as rubidium and cesium) or to the alkali-earth group (such as strontium) through optical tweezers [9, 12]. The atoms, which encode the physical qubits are then placed in a graph structure (current technology allows to place atoms on a 2D lattice), called quantum register. The qubit basis states denoted as $|0\rangle$ and $|1\rangle$ can be encoded on different pairs of electronic states of the atom, depending on the usage mode, i.e. digital or analog. From now on, we will consider rubidium as it is the atomic species adopted by both QuEra and Pasqal companies, leading start-ups in the field of quantum computation with neutral atoms in USA and Europe, respectively. In digital mode, the basis states $\{|0\rangle,|1\rangle\}$ are encoded in the hyperfine structure levels of the electronic state $|5S_{1/2}\rangle$ with $F=1$ and $F=2$, respectively (see fig. 1 in [12]). In contrast, in the analog mode, the state $|0\rangle$ is encoded in the $|g\rangle=|5S_{1/2}\rangle$ ground state of rubidium, while the state $|1\rangle$ corresponds to the highly excited Rydberg level $|r\rangle=|nS_{1/2}\rangle$ with $50\leq n\leq 70$ [9, 12].
@@
The system dynamics is driven by means of laser pulses with a tunable Rabi frequency $\Omega$ (proportional to the laser amplitude), detuning frequency $\delta$ (i.e. the difference between the qubit resonance frequency and the actual laser frequency) and their relative phase $\phi$. The crucial key ingredient of quantum computing with neutral atoms are the strong Van der Waals interactions that are experimented by neighboring atoms in the excited Rydberg state [25, 26, 27], which hinder the simultaneous excitation of atoms whose distance is smaller than a threshold length known as the Rydberg blockade radius [28]. This is due to a shift up in energy of the $|rr\rangle$ of one of the two atoms caused by the interaction term [25] which physically decouples such state form the dynamics. As a consequence, neutral atoms represent a very powerful setup to achieve high-fidelity entanglement [29] since whenever a couple of qubits is placed within the blockade radius, the attempt to excite them would result in the final entangled state $\frac{|gr\rangle+|rg\rangle}{\sqrt{2}}$. Appendix from A.1 to A.4 illustrate the details of digital and analog computational modes together with register loading and readout technicalities.
@@ 3.3.1 Gate fidelity improvement
The future perspective for neutral atoms devices points towards the possibility to implement digital gates with local addressing laser pulses. This requires a considerable hardware improvement, consisting mainly in the massive development of the current capability to address single atoms with lasers. In addition, the correction of the main errors, which hinder a perfect implementation of quantum gates, has to be taken into account. Here we look at the strategies adopted to attain the highest possible fidelity for 2- and 3- qubit gates, typically the CZ (involved in the preparation of Bell pairs) and the CCZ, useful for more complex quantum algorithms.
@@ 4 Graph problems
Graph problems are of crucial importance due to their intrinsic potential to revolutionize our society in many research fields. The development of quantum computers is expected to boost industrial processes [16], chemistry and medicine, leveraging machine learning protocols [81, 82], with a special focus on drug discovery [83], molecular docking [84], and many body physics [17].
@@ 4.1 MIS, Max-Cut and Max-Clique
Notably, this Hamiltonian is equivalent to that in eq. (10). As a consequence, the MIS is a native problem solvable on a neutral atom device since the ground state of the Hamiltonian (10), which can be properly prepared, encodes the solution of the MIS. the more general Maximum Weight Independent Set (MWIS) problem, where edges in $E$ are given a weight, is more complicated and, from an hardware point of view, it can be handled by encoding the weights into the strength of the interactions between atoms, i.e. choosing a different distance between the atoms. Let us notice that not all possible graphs and sets of weight can be implemented in an exact way, a problem that will be addressed in section 4.3.
@@
With some manipulations of the graph, both the MaxCut and the MaxClique can be mapped into a MIS [80]. In particular, the MaxClique is equivalent to a MIS on the complementary graph of $G$. As a consequence, if a given problem can be written as either a MIS, a MaxCut or a MaxClique, it can be directly implemented and solved with neutral atom arrays. For completeness, we also mention that clustering aggregation problems [85] can be restated in terms of a MIS.
To the best of our knowledge, the first proof of the possibility to solve the MIS on a neutral atoms device dates back to 2018 [86], working with a pretty straightforward implementation and the idea of improving available algorithms [87, 88]. The developments pointed toward two main directions: the improvement of hybrid and variational algorithms and the development of techniques for mapping the graphs on the qubits register. This approaches will be discussed in sec. 4.2 and 4.3, respectively.
@@ 4.2 QAOA and QAA for MIS
Although a standard version of QAOA is enough to find good approximate solutions, even in shallow circuit regimes and low repetition rate, especially if coupled to a bayesian optimizer [89], a direct comparison against its main competitor, i.e. the Variational Quantum Adiabatic Algorithm (VQAA) which is an improved version of QAA where the adiabatic pulse is optimized classically, emphasizes how the QAOA can be outperformed by VQAA even at small effective depth [90], at least for MIS instances up to 289 qubits. In particular, Ebadi et al. showed that hardware limitations, including leakage errors, blockade and pulse imperfections, hinder the classical optimizer to spot the best QAOA parameters and that’s why, when increasing the circuit depth $p$, the approximation ratio reaches a maximum and then starts to fall again. The same happens for VQAA but the peak of the approximation ratio is much higher than for QAOA, reaching a maximum value greater than $90\%$ [90]. Furthermore, VQAA has also been benchmarked with Simulated Annealing (SA) [91]. The approximation ratio for SA at low depth is similar to that of VQAA. Nonetheless, for some graph instances with many local minima (corresponding to independent sets with one vertex less than the MIS) SA gets stuck and the approximation ratio displays a plateau until a very high depth is reached [90]). The performance of the VQAA is described by quasi-adiabatic evolution in agreement with the Landau-Zener formula [92], which qualitatively indicates that the fidelity, i.e. the probability to measure a MIS solution bitstring, strongly depends on the energy gap between the ground state (MIS energy) and the first excited state: the higher the latter, the higher the fidelity [90].
@@
The former consists of a QAOA-like quantum evolution which does not require further classical optimization. The underlying idea is to achieve a monotonic decrease of the expectation value of the problem Hamiltonian over layers. This is done by computing the evolution parameter $\beta_{j}$ of the mixer layer $j$ from the expectation value of the commutator between the cost and the mixer Hamiltonians in the previous layer, i.e. $\beta_{j}=-k\langle\psi_{j-1}|i[H_{M},H_{C}]|\psi_{j-1}\rangle$. By choosing $k=2$, which is enough to achieve $\frac{d}{dt}\langle H_{C}\rangle\leq 0$ according to the authors [93], a continuous decrease in energy $\langle H_{C}\rangle$ is induced. As a consequence, an approximate solution of the ground state naturally emerges at the end of the circuit. A benchmark of MaxCut graph instances with $n=2,3$ sites is provided, showing that the approximation ration approaches $1$ as the number of layers increases. Unfortunately, larger system sizes are not taken into account.
@@ 4.3 Quantum wiring and problem mappings
By construction, the quantum register containing Rydberg atoms has limited connectivity, equivalent to that of a Unitary Disk (UD) graph. It is therefore fundamental to figure out how is it possible to address those graph problems which are natively non-UD. Quantum wires are of great help in this sense [96]. This approach consists in connecting far nodes with a chain of atoms of even length (to satisfy the MIS constraints), as shown in [96]. An alternative arrangement of the wires relies on orthogonal grid drawing (ODG) [97] and it is presented in [98]. The main difference consists of the positions of the ancillary qubits, which are displaced in an orthogonal pattern [98]. The number of ancillary qubits scales sublinearly in the number of qubits of the original problem. An inverse mapping is provided in [99], where the MIS has been solved on platonic solids mapped to 2D surfaces.
@@ 6.1.1 Ising model ground state in the Ground-Rydberg basis
Far from being a trivial and effortless task, the preparation of the ground state of the Ising model with high fidelity, with a focus on the antiferromagnetic (J=1J=-1) version, has attracted a lot of attention since the dawn of quantum computation with Rydberg atoms [124, 125, 126, 127, 128]. A standard annealing scheme, in the Ground-Rydberg encoding basis, is typically employed to accomplish this goal. Basically, it works as follows:
@@ 7.1 Molecular Docking and Drug Discovery
The design of new drugs has always been a crucial task in modern medicine [174]. In recent years, new computational frameworks and techniques have been exploited to this aim, including machine learning [83, 177] and, ultimately, quantum computing [178, 179]. The steps of the process leading to the discovery of new drugs, as illustrated in [84], include the solution of a molecular docking problem [180, 84, 181]. The latter consists in finding the best configuration of a molecule (the ligand) such that it perfectly ties with the binding site of a protein (the target). Each possible configuration found by the algorithm is evaluated with a score. More details on the ranking metric of protein-ligand configurations and Hamiltonian reformulation of the problem can be found in [182]. To tackle a molecular docking problem on a neutral atoms device, it is required to map it into a Maximum Weighted Independent Set (MWIS), as explained in [183]. In this paper, Garrigues et al. compared the performance of Variational Quantum Adiabatic Algorithm (VQAA), which makes use of a modified version of the Bayesian optimizer [89] called Hyperopt optimizer [184], in order to optimize the hyperparameters (rise, sweep and fall time together with the detuning and the Rabi frequencies), with that of a supervised machine learning version of QAA, dubbed MLQAA, which predicts the best QAA parameters based on a training set of given VQAA runs. Each algorithm is evaluated by means of a scoring function, ranging between 0 and 1, which takes into account both energetic and geometric factors, evaluating how well the ligand fits the receptor’s binding site and interaction energies between the ligand and the receptor. Because of the high variability of optimal QAA parameters, MLQAA average score is about one half of the one achieved by VQAA [183]. Nevertheless, the score achieved by VQAA is almost 0.8 after 500 iterations and seems to be independent of the number of nodes (up to 11 at least). If this is going to be further confirmed, this approach may suggest potential scalability advantages on larger problem sizes.

### v2

@@ 1 Introduction
The report is structured as follows. Sec. 2 introduces the basic principles of quantum computing with neutral atom QPUs, whose performances are benchmarked in Sec. 3. Sec. 4 deals with classical optimization problems on graphs, focusing on the implementation of variational quantum algorithms to solve them. Sec. 6 describes applications directly related to Physics, especially to quantum many-body theory, while applications to the fields of chemistry and pharmacology are presented in Sec. 7. Sec. 8 is devoted to machine learning enhancement with the help of Rydberg atom arrays. Finally, the description of some other relevant problems, such as the SAT [176] and QUBO [126] ones, is provided in Sec. 5. We draw some conclusions in Sec. 9, providing some details on the hardware in the Appendix.The literature on Quantum Computing and Simulations has surged in recent years, because these cutting-edge paradigms of computation are expected to bring disruptive improvements in our current computational capacity, enhancing the algorithmic performance in solving a huge variety of problems of interest for scientific and industrial applications. A variety of algorithmic methods have been proposed to tackle complex classical and quantum problems in both the digital [85] and the analog approach [77].
Most of the application cases discussed in this review, such as QAOA and QAA-based optimization, quantum chemistry via VQE, and the simulation of lattice models, can be realized in principle on other quantum platforms, among which are superconducting qubits, trapped ions, photonics, and NMR systems. However, the practical feasibility of each of these implementations depends fundamentally on the native connectivity and coherence properties of the hardware, along with its control capabilities. In this respect, neutral-atom processors really offer singular advantages for problems whose natural form is a geometric or graph-constrained Hamiltonian, where the Rydberg interaction directly encodes the underlying combinatorial structure.
@@ 2 Quantum Computing with neutral atoms
Figure 1: (a) Atomic levels of Rubidium used for computation: the ground state $|5S_{1/2}\rangle$, the intermediate state $|6P_{3/2}\rangle$ and the Rydberg level $|nS_{1/2}\rangle$ where $50\leq n\leq 90$ (here we choose $n=60$). In the digital approach, the computational basis is formed by the hyperfine ground state with total angular momentum $F=1$ and $F=2$. In the analog approach the computational basis is given by the ground and the Rybderg level, and the hyperfine structure is neglected for the purpose of the computation. The atom can be excited to the Rydberg level via a two-photon transition scheme, including a first excitation to the intermediate level with a 420nm420\;nm wavelength photon and a subsequent excitation to the Rydberg level via a 1013nm1013\;nm photon. (b) Rydberg blockade effect: when two atoms are at a distance $R$, the doubly excited level $|rr\rangle$ is shifted up in energy of a quantity $U=C_{6}/R^{6}$.
@@
A neutral atom Quantum Processing Unit (QPU) works by trapping and cooling atoms, which typically belong to the alkali group (such as rubidium and cesium) or to the alkali-earth group (such as strontium) through optical tweezers [89, 225]. The atoms, which encode the physical qubits are then placed in a graph structure (current technology allows to place atoms on a 2D lattice), called quantum register. The qubit basis states denoted as $|0\rangle$ and $|1\rangle$ can be encoded on different pairs of electronic states of the atom, depending on the usage mode, i.e. digital or analog. From now on, we will consider rubidium as it is the atomic species adopted by both QuEra and Pasqal companies, leading start-ups in the field of quantum computation with neutral atoms in USA and Europe, respectively. In digital mode, the basis states $\{|0\rangle,|1\rangle\}$ are encoded in the hyperfine structure levels of the electronic state $|5S_{1/2}\rangle$ with $F=1$ and $F=2$, respectively. In contrast, in the analog mode, the state $|0\rangle$ is encoded in the $|g\rangle=|5S_{1/2}\rangle$ ground state of rubidium, while the state $|1\rangle$ corresponds to the highly excited Rydberg level $|r\rangle=|nS_{1/2}\rangle$ with $50\leq n\leq 90$ [89, 225]. Fig. 1(a) shows a simple scheme of the electronic levels of Rubidium needed for computation.
@@
The system dynamics is driven by means of laser pulses with a tunable Rabi frequency $\Omega$ (proportional to the laser amplitude), detuning frequency $\delta$ (i.e. the difference between the qubit resonance frequency and the actual laser frequency) and their relative phase $\phi$. The crucial key ingredient of quantum computing with neutral atoms are the strong Van der Waals interactions that are experimented by neighboring atoms in the excited Rydberg state [224, 145, 1], which hinder the simultaneous excitation of atoms whose distance is smaller than a threshold length known as the Rydberg blockade radius [207]. This is due to a shift up in energy of the $|rr\rangle$ of one of the two atoms caused by the interaction term [224] which physically decouples such state form the dynamics (see fig. 1(b)). As a consequence, neutral atoms represent a very powerful setup to achieve high-fidelity entanglement [100] since whenever a couple of qubits is placed within the blockade radius, the attempt to excite them would result in the final entangled state $\frac{|gr\rangle+|rg\rangle}{\sqrt{2}}$. Appendix from A.1 to A.4 illustrate the details of digital and analog computational modes together with register loading and readout technicalities. Neutral-atom platforms also exhibit a number of unique capabilities that set them apart from other quantum technologies. First, programmable all-to-all connectivity through dynamical rearrangement of optical tweezers allows the physical geometry of the qubit register to be adapted to the structure of the problem. Second, Rydberg-mediated interactions, such as the blockade mechanism, naturally implement hard constraints which maps directly onto the Hamiltonians of combinatorial problems. Third, the large scalability potential, arising from trapping hundreds of atoms in parallel, provide a versatile computational framework.
@@ 3.1 Register mapping improvement
Figure 2: Workflow of Geyser [159] and Graphine [160], whose details are described in the main text.
@@
Fig. 2 schematically illustrates the workflow of both Geyser and Graphine.
@@ 3.3 Gates implementation and Quantum Error Correction
CZ CCZ
Input Output Input Output
$\rangle |00|00\rangle$ $\rangle |000|000\rangle$
$\rangle |001|001\rangle$
$\rangle |01|01\rangle$ $\rangle |010|010\rangle$
$\rangle |011|011\rangle$
$\rangle |10|10\rangle$ $\rangle |100|100\rangle$
$\rangle |101|101\rangle$
$\rangle |11-|11\rangle$ $\rangle |110|110\rangle$
$\rangle |111-|111\rangle$
Table 1: Effect of quantum gates CZ and CCZ on computational basis states.
@@ 3.3.1 Gate fidelity improvement
The future perspective for neutral atoms devices points towards the possibility to implement digital gates with local addressing laser pulses. This requires a considerable hardware improvement, consisting mainly in the massive development of the current capability to address single atoms with lasers. In addition, the correction of the main errors, which hinder a perfect implementation of quantum gates, has to be taken into account. Here we look at the strategies adopted to attain the highest possible fidelity for 2- and 3- qubit gates, typically the CZ (involved in the preparation of Bell pairs) and the CCZ, useful for more complex quantum algorithms. The effect of quantum gates CZ and CCZ on computational basis states are shown, for the sake of clarity, in Tab. 1.
@@ 4 Graph problems
Graph problems are of crucial importance due to their intrinsic potential to revolutionize our society in many research fields and the growing availability of advanced compilation frameworks for optimization on Rydberg hardware [184]. The development of quantum computers is expected to boost industrial processes [226], chemistry and medicine, leveraging machine learning protocols [96, 35], with a special focus on drug discovery [74], molecular docking [66], and many body physics [17].
@@ 4.1 MIS, Max-Cut and Max-Clique
Figure 3: MIS, MaxCut and MaxClique solution for a simple graph instance
@@
Notably, this Hamiltonian is equivalent to that in eq. (10). As a consequence, the MIS is a native problem solvable on a neutral atom device since the ground state of the Hamiltonian (10), which can be properly prepared, encodes the solution of the MIS. the more general Maximum Weight Independent Set (MWIS) problem, where edges in $E$ are given a weight, is more complicated and, from an hardware point of view, it can be handled by encoding the weights into the strength of the interactions between atoms, i.e. choosing a different distance between the atoms. Let us notice that not all possible graphs and sets of weight can be implemented in an exact way, as also highlighted in recent studies addressing hardware compatibility and embedding overhead on Rydberg arrays [184], a problem that will be addressed in section 4.3.
@@
With some manipulations of the graph, both the MaxCut and the MaxClique can be mapped into a MIS [157]. In particular, the MaxClique is equivalent to a MIS on the complementary graph of $G$. As a consequence, if a given problem can be written as either a MIS, a MaxCut or a MaxClique, it can be directly implemented and solved with neutral atom arrays. For completeness, we also mention that clustering aggregation problems [127] can be restated in terms of a MIS. Fig. 3 displays the solution for the three aforementioned graph problems on a simple 7-node instance. This graph has been properly chosen such that the solution node set is different for each problem.
To the best of our knowledge, the first proof of the possibility to solve the MIS on a neutral atoms device dates back to 2018 [165], working with a pretty straightforward implementation and the idea of improving available algorithms [49, 52]. The developments pointed toward two main directions: the improvement of hybrid and variational algorithms and the development of techniques for mapping the graphs on the qubits register. Recent works have further advanced these directions by integrating quantum-informed reduction strategies [185] and by developing comprehensive compilation pipelines for mapping generic MIS instances onto Rydberg hardware [184]. This approaches will be discussed in sec. 4.2 and 4.3, respectively.
@@ 4.2 QAOA and QAA for MIS
Although a standard version of QAOA is enough to find good approximate solutions, even in shallow circuit regimes and low repetition rate, especially if coupled to a bayesian optimizer [204], a direct comparison against its main competitor, i.e. the Variational Quantum Adiabatic Algorithm (VQAA) which is an improved version of QAA where the adiabatic pulse is optimized classically, emphasizes how the QAOA can be outperformed by VQAA even at small effective depth [63], at least for MIS instances up to 289 qubits. In particular, Ebadi et al. showed that hardware limitations, including leakage errors, blockade and pulse imperfections, hinder the classical optimizer to spot the best QAOA parameters and that’s why, when increasing the circuit depth $p$, the approximation ratio reaches a maximum and then starts to fall again. The same happens for VQAA but the peak of the approximation ratio is much higher than for QAOA, reaching a maximum value greater than $90\%$ [63]. Complementary approaches have demonstrated that Bayesian Optimization can be used to design highly efficient quantum annealing schedules, achieving substantial improvements in fidelity for MIS instances on Rydberg devices [69]. Furthermore, VQAA has also been benchmarked with Simulated Annealing (SA) [209]. The approximation ratio for SA at low depth is similar to that of VQAA. Nonetheless, for some graph instances with many local minima (corresponding to independent sets with one vertex less than the MIS) SA gets stuck and the approximation ratio displays a plateau until a very high depth is reached [63]). The performance of the VQAA is described by quasi-adiabatic evolution in agreement with the Landau-Zener formula [87], which qualitatively indicates that the fidelity, i.e. the probability to measure a MIS solution bitstring, strongly depends on the energy gap between the ground state (MIS energy) and the first excited state: the higher the latter, the higher the fidelity [63]. Furthermore, recent analytical studies have shown that the time-to-solution for MIS instances on Rydberg devices is intrinsically constrained by graph-dependent hardness parameters and degeneracy structure, setting fundamental performance limits for adiabatic and annealing-based approaches [5].
@@
The former consists of a QAOA-like quantum evolution which does not require further classical optimization. The underlying idea is to achieve a monotonic decrease of the expectation value of the problem Hamiltonian over layers. This is done by computing the evolution parameter $\beta_{j}$ of the mixer layer $j$ from the expectation value of the commutator between the cost and the mixer Hamiltonians in the previous layer, i.e. $\beta_{j}=-k\langle\psi_{j-1}|i[H_{M},H_{C}]|\psi_{j-1}\rangle$. By choosing $k=2$, which is enough to achieve $\frac{d}{dt}\langle H_{C}\rangle\leq 0$ according to the authors [128], a continuous decrease in energy $\langle H_{C}\rangle$ is induced. As a consequence, an approximate solution of the ground state naturally emerges at the end of the circuit. A benchmark of MaxCut graph instances with $n=2,3$ sites is provided, showing that the approximation ratio approaches $1$ as the number of layers increases. Unfortunately, larger system sizes are not taken into account.
@@ 4.3 Quantum wiring and problem mappings
Figure 4: Example of quantum wiring on a 8-node atom arrangement. Nodes 11 and 66 are connected through a wire of an even number of atoms to ensure MIS constraint. Quantum wiring unlocks the possibility to solve the MIS problem on a non-UD graph by connecting far nodes with chains of extra atoms
By construction, the quantum register containing Rydberg atoms has limited connectivity, equivalent to that of a Unitary Disk (UD) graph. It is therefore fundamental to figure out how is it possible to address those graph problems which are natively non-UD. Quantum wires are of great help in this sense [110]. This approach consists in connecting far nodes with a chain of atoms of even length (to satisfy the MIS constraints), as shown in Fig. 4. An alternative arrangement of the wires relies on orthogonal grid drawing (ODG) [62] and it is presented in [51]. The main difference consists of the positions of the ancillary qubits, which are displaced in an orthogonal pattern [51]. The number of ancillary qubits scales sublinearly in the number of qubits of the original problem. An inverse mapping is provided in [29], where the MIS has been solved on platonic solids mapped to 2D surfaces.
@@ 6.1.1 Ising model ground state in the Ground-Rydberg basis
Figure 5: Adiabatic sequence to prepare the ground state of the antiferromagnetic Ising model. The algorithm consists of three main phases called "rise", "sweep" and "fall": starting from a far negative detuning $\delta$ and a null Rabi frequency $\Omega$, the latter is raised up to a maximum value $\Omega_{max}$; after that, while keeping $\Omega=\Omega_{max}$, the detuning is brought to a far positive value $\delta_{max}$ and, at the very end, the Rabi frequency is turned off.
Far from being a trivial and effortless task, the preparation of the ground state of the Ising model with high fidelity, with a focus on the antiferromagnetic (J=1J=-1) version, has attracted a lot of attention since the dawn of quantum computation with Rydberg atoms [117, 118, 218, 56, 177]. A standard annealing scheme, in the Ground-Rydberg encoding basis, is typically employed to accomplish this goal. The full sequence is provided in Fig. 5. Basically, it works as follows:
@@ 7.1 Molecular Docking and Drug Discovery
The design of new drugs has always been a crucial task in modern medicine [15]. In recent years, new computational frameworks and techniques have been exploited to this aim, including machine learning [74, 11] and, ultimately, quantum computing [212, 168]. The steps of the process leading to the discovery of new drugs, as illustrated in [66], include the solution of a molecular docking problem [146, 66, 193]. The latter consists in finding the best configuration of a molecule (the ligand) such that it perfectly ties with the binding site of a protein (the target) [112]. Each possible configuration found by the algorithm is evaluated with a score. To tackle a molecular docking problem on a neutral atoms device, it is required to map it into a Maximum Weighted Independent Set (MWIS), as explained in [73]. In this paper, Garrigues et al. compared the performance of Variational Quantum Adiabatic Algorithm (VQAA), which makes use of a modified version of the Bayesian optimizer [204] called Hyperopt optimizer [16], in order to optimize the hyperparameters (rise, sweep and fall time together with the detuning and the Rabi frequencies), with that of a supervised machine learning version of QAA, dubbed MLQAA, which predicts the best QAA parameters based on a training set of given VQAA runs. Each algorithm is evaluated by means of a scoring function, ranging between 0 and 1, which takes into account both energetic and geometric factors, evaluating how well the ligand fits the receptor’s binding site and interaction energies between the ligand and the receptor. Because of the high variability of optimal QAA parameters, MLQAA average score is about one half of the one achieved by VQAA [73]. Nevertheless, the score achieved by VQAA is almost 0.8 after 500 iterations and seems to be independent of the number of nodes (up to 11 at least). If this is going to be further confirmed, this approach may suggest potential scalability advantages on larger problem sizes.